OpenStudy (anonymous):

Choose the missing term in the following pair of equivalent ratios. 50 : 9 = ___ : 27 75 100 125 150

OpenStudy (jango_in_dtown):

suppose the missing term is x. Then 50/9=x/27 this gives x=(50*27/9)=150

OpenStudy (anonymous):

@texaschic101 can u help me?

OpenStudy (texaschic101):

ratios are nothing but equivalent fractions..50:9 is the same as 50/9. 50/9 = x/27.....to find an equivalent fraction, you need to multiply numerator and denominator by the same number...notice how the 9, when multiplied by 3 is 27...and since you multiplied the denominator by 3, you have to multiply the numerator by 3. So what is 50 * 3 ?

OpenStudy (anonymous):

150

OpenStudy (texaschic101):

or just solve it like a proportion.. 50/9 = x/27..cross multiply, then solve for x yes... 50/9 = 150/27

OpenStudy (anonymous):

im closing the qeustion now

OpenStudy (texaschic101):

ok :)

OpenStudy (anonymous):

texaschic can u help me with algebra?

OpenStudy (texaschic101):

possibly....do you have an open question ?